Pleas help me choose one of the answers SHOWN Question 7 (10 points) Saved Calculate the equivalent annual cost for years 1 through infinity of $1,000,000 now plus $500,00
Answer #1

Answer
A=PV/i
PV=amount now + PV of amount 5 years from now
PV=FV*(1+i)^(-n)
FV=500000
i=interest rate =5%
n=years=5
A=annual cost
PV=500000*(1.05^-5)+1000000
=1391763.08
A=PV*i=1391763.08*0.05
=69588.154
=69763 (approximately)

Option 5